4. A rubber band of dimensions 10 mm x 2 mmx 200 m
Answer #1

Dimensions of rubber band = 10 mm x 2 mm x 200 mm.

Volume of rubber band = 4000 mm3

= 4.0x10-6 m3

Temperature, T = 20 oC = 293 K

Chains / unit m3 = 1.9x1026

Work done per unit volume, w = -1/2 \nukT (\lambda2 + 2/\lambda - 3)

= -1/2 * 1.9x1026 * 1.38x10-23 * 293 (52 + 2/5 - 3)

= -1/2 * 1.9x1026 * 1.38x10-23 * 293 * 22.4

= - 4.51x106 J / m3

Total work done, W = - 4.51x106 J / m3 * 4.0x10-6 m3

= - 18.04 J

\DeltaU = \DeltaW + \DeltaQ

Since, \DeltaU = 0

\DeltaQ = - \DeltaW

\DeltaQ = 18.04 J

So, Heat lost = \DeltaQ = 18.04 J